OpenStudy (anonymous):

Let g(x) = x2 + x. Find g(-10). does that mean i put 10 for all the x's?

OpenStudy (dan815):

syes

OpenStudy (anonymous):

yes,

OpenStudy (dan815):

g(-10) = 90

OpenStudy (anonymous):

my bad, dan815 is right

jimthompson5910 (jim_thompson5910):

you replace all copies of x with -10 and not 10

jimthompson5910 (jim_thompson5910):

g(x) = x^2 + x g(-10) = (-10)^2 + (-10) g(-10) = 100 - 10 g(-10) = 90

OpenStudy (anonymous):

oh ok thank you!!

OpenStudy (dan815):

jimthompson medal stealer!!

OpenStudy (anonymous):

ummm?

jimthompson5910 (jim_thompson5910):

there you go, it's just a silly medal...

OpenStudy (dan815):

YAY :D